OpenStudy (anonymous):

Find the equation in x and y for the line tangent to the curve x(t)=2/t, y(t)=t^2+2 at the point (1,6).

OpenStudy (anonymous):

I know that you usually take the derivative of an x or y if you have it, but I don't know how to combine them both in a single equation.

OpenStudy (mathmale):

You'll need to find the derivative, dy/dx, as usual. In this case y ou're working with "parametric equations," where t is your parameter.

OpenStudy (mathmale):

Your dy/dx is\[\frac{ dy }{ dx }=\frac{ \frac{ dy }{ dt } }{ dx/dt }\]

OpenStudy (mathmale):

I'd guess you already know how to find dy/dt and dx/dt.

OpenStudy (anonymous):

x'(t)=4, y'(t)=4t^3.

OpenStudy (anonymous):

WOAH wait wrong set

OpenStudy (mathmale):

You'll have to take the given point (1,6) and work backward to find the t value that gives you x and y.

OpenStudy (anonymous):

x'(t)=-2/x^2, y'(t)=2t

OpenStudy (mathmale):

If x(t)=t/2, find dx/dt. It's not 4. If y(t)=t^2+2, find dy/dt.

OpenStudy (anonymous):

My bad, it's x(t)=2/t.

OpenStudy (mathmale):

Well, if x(t)=2/t, then x(t) = 2*(1/t) or 2*t^(-1). Find dx/dt.

OpenStudy (mathmale):

I'd use the power rule, were I you.

OpenStudy (mathmale):

@RightInTheCranium: sorry, but I have to ask you to hurry. Bedtime for me! ;)

OpenStudy (anonymous):

Er...x'(t)=-2/t^2 and y'(t)=2t isn't right?

OpenStudy (mathmale):

If x(t)=2/t=2t^(-1), then x '(t) = -2/t^2, yes. Good. Now we're on the right track. Go back and look at my formula for finding dy/dx as a function of dy/dt and dx/dt. Write out your dy/dx for this particular problem.

OpenStudy (anonymous):

\[\frac{ 2t }{-2/t^2 }\]?

OpenStudy (mathmale):

\[\frac{ dy }{ dx }=\frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } }=?\]

OpenStudy (mathmale):

Good. Now reduce that result of yours.

OpenStudy (anonymous):

Bleh, typing in this site can be so slow sometimes. It's -t^3.

OpenStudy (mathmale):

Yes. Now we have to find the value of t so that we can find the value of this derivative.

OpenStudy (mathmale):

We're given a point on the curve; it is (1,6). what is the formula for x(t)? Type it here.

OpenStudy (anonymous):

x(t)=2/t

OpenStudy (mathmale):

Good. Now, the x value of the given point is 1. Set y our formula for x(t) equal to 1 and solve the resulting equation for t.

OpenStudy (mathmale):

t = ?

OpenStudy (anonymous):

t=2. So -t^3=-8 ?

OpenStudy (mathmale):

Yes. That's the slope of the tangent line to the curve at (1,6). This tangent line obviously goes thru (1,6). Use your slope and this point to write the equation of the tangent line.

OpenStudy (anonymous):

y-6=-8(x-1) => y-8x+2=0 ?

OpenStudy (anonymous):

Or wait

OpenStudy (anonymous):

y-8x-14=0

jimthompson5910 (jim_thompson5910):

Alternative Route to find the slope: solve x = 2/t for t to get t = 2/x Then plug that into y = t^2 + 2 to get y = (4/x^2) + 2 now apply the derivative. I'm skipping a bunch of steps, but you'll end up with dy/dx = -8/x^3 Plugging in x = 1 leads to dy/dx = -8 The first method is more effective because often in parametric mode, it's hard to isolate t.

OpenStudy (mathmale):

Your y-6=-8(x-1) is correct. You could put it into a fancier form later if you wish.

OpenStudy (anonymous):

Hmm...I'm looking at the answer choices I have for this questions and none of them look like what I got.

OpenStudy (mathmale):

I'd suggest you double check to ensure that you have your parametric equations for x and y correct.

OpenStudy (anonymous):

Seems alright to me, but obviously there's something not quite right. Bloody hell. Oh well, time to ring up the professor. Thank you anyway!

OpenStudy (mathmale):

Take care. Can't resist saying I hope you'll try this problem again on your own before seeing (or ringing up) your prof. We've covered a lot of ground and made a lot of progress.

jimthompson5910 (jim_thompson5910):

what are your answer choices @RightInTheCranium ?

OpenStudy (mathmale):

Signing off now. Good night. Jim: Thanks for your input!

OpenStudy (anonymous):

\[2x+\frac{ 5 }{ 4 }\] \[4x-\frac{ 11 }{ 2 }+\frac{ y }{ 4 }\] \[2x-\frac{ 4 }{ 4 }\] \[4x-7+\frac{ y }{ 2 }\] -4x+3

jimthompson5910 (jim_thompson5910):

hmm that's strange, none of them are equivalent to y = -8x+14. I was hoping at least one was

OpenStudy (anonymous):

Hmm...I solved for t from the x(t) equation. Maybe the y(t) will yield something.

OpenStudy (anonymous):

Hm, gonna keep trying. Setting t^2+2=6 might give me something. Hopefully?

jimthompson5910 (jim_thompson5910):

I already posted that above in a previous post

OpenStudy (anonymous):

JUST when you think it's going to be easy. I'm don't know what I expected. Ugh. Time to consult the professor. Thanks anyway! (Again.)

OpenStudy (triciaal):

I got the last option here is why i would choose that

OpenStudy (triciaal):